OpenStudy (anonymous):

Solve. 64^x= 16^(x–1)

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\large 64^x = 16^{x-1} \implies 4^{3x} = 4^{2(x-1)} \implies 4^{3x} = 4^{2x-2}\] Just equate the exponents or powers and find x..

OpenStudy (anonymous):

can you rewrite the 64 and 16 as powers with base 2?

OpenStudy (anonymous):

or do as in waterin eyes case, base 4....

OpenStudy (anonymous):

What do you mean equate the exponents or powers?

OpenStudy (anonymous):

What are the exponents tell me on LHS as well as on RHS first tell me this..

OpenStudy (maheshmeghwal9):

Rule: -\[\huge{a^x=a^y \implies x=y.}\]

OpenStudy (anonymous):

\[\huge (x)^a = (x)^b \implies \color{blue}{a = b}\]

OpenStudy (anonymous):

So just equate the coefficients there.. Equate means just make them equal or between them use "=" and then solve for x.

OpenStudy (anonymous):

*exponents not coefficient..

OpenStudy (anonymous):

So 3x=2x-2?

OpenStudy (anonymous):

Go ahead..

OpenStudy (anonymous):

Yes you are right... @HayleyWhite17

OpenStudy (anonymous):

6x-6?

OpenStudy (anonymous):

How ??

OpenStudy (anonymous):

\(3x = 2x - 2\) Subtract 2x both the sides and tell me what do you get ??

OpenStudy (anonymous):

Ok hold on